# ReferencesSuppReferences

### v3

@@ 2.2.1 Potential Outcomes Framework
We assume that $d\in\mathbf{D}$ for an appropriate set $\mathbf{D}$. A potential outcome $Y_{t}\left(d\right)$ describes which value the outcome would have taken at time $t$ under treatment value $d$. The definition implies that the potential outcome $Y_{t}\left(d\right)$ does not depend on future treatments. Rambachan and Shephard (2021) described this property as “non-anticipating potential outcomes”. The definition also implies that the potential outcome $Y_{t}\left(d\right)$ does not depend on past treatments (unless they are elements of $Z_{t}$). This is realistic in the high-frequency event study setting for two reasons. First, $Y_{t}$ typically measures a change in an asset price or a survey forecast within a narrow window around a policy announcement. Thus, market efficiency or rational expectations, respectively, imply that market participants or professional forecasters use all public information available at the start of the window and so $Y_{t}$ is constructed by conditioning on that information. Second, for $t\in\mathbf{P}$ the latest past treatment is $D_{t-1}$ ($(\{t-1\}\in\mathbf{C})$ which captures the change in a policy variable within a 30-minute window that does not involve any policy announcement and so the change is very small or equal zero. The potential outcome should not be confused with the outcome $Y_{t}=Y_{t}\left(D_{t}\right)$. Finally, note that in terms of the structural function in (2.3), $Y_{t}(d)=\varphi_{Y}(d\,,Z_{t}\,,u_{t}\,,t)$.
@@
Under these assumptions, the MCE of the time $t$ policy on the time $t$ normalized potential outcome is defined as $\partial\widetilde{Y}_{t}\left(d\right)/\partial d$. Assumption 5-6 involve the normalized quantities instead of the actual quantities since we analyze estimands and estimators under relative exogeneity (2.4), for which the support of $e_{t}$ and $D_{t}$ necessarily become unbounded. Following Rambachan and Shephard (2021), to apply basic tools such as the fundamental theorem of calculus we need the argument of the relevant function to have support on a closed interval.
@@ 2.3 Identification Results
where $H_{t}\left(d\right)=\mathbf{1}\{d\leq\widetilde{D}_{t}\}(\widetilde{D}_{t}-\mathbb{E}(\widetilde{D}_{t}))$ with $\mathbb{E}(H_{t}\left(d\right))\geq 0,$ $\int_{\mathbf{D}}\mathbb{E}\left(H_{t}\left(d\right)\right)\mathrm{d}d=1$ and
$\Delta_{t}=\mathbb{E}\left[\widetilde{Y}_{t}\left(\underline{d}\right)\left(\widetilde{D}_{t}-\mathbb{E}\left(\widetilde{D}_{t}\right)\right)\right]$
is decreasing in $\sigma_{e,t}^{2}$.
@@
For the weighted average of the MCEs in Theorem 2, a higher weight is not necessarily given to large values of $d$ because large values of $d$ may be associated with small tail probabilities of the distribution of $\widetilde{D}_{t}.$ The intuition underlying how $\beta_{\mathrm{ES},t}$ is able to recover a weighted average of MCEs under relative exogeneity is as follows. Relative exogeneity, in combination with Assumptions 1-3, imply that when appropriately normalized, the potential outcome $Y_{t}\left(d\right)$ and the policy variable $D_{t}$ behave as if they are uncorrelated. Intuitively, the effect of the policy shock on the policy variable is not influenced by $Y_{t}$ (separability) and so changes in the policy variable are entirely determined by changes in the policy shock (relative exogeneity). In addition, as $\sigma_{e,t}^{2}\rightarrow\infty$ the relative bias generated by the omitted variables $Z_{t}$ becomes negligible since the correlation between $Z_{t}$ and $D_{t}$ is an order of magnitude smaller than the variation in $D_{t}$ generated by $e_{t}$.
@@
The key identification condition is relative exogeneity, which also relates to the size of the time window surrounding an announcement or event. As the size of the window expands, it becomes less likely that the policy shock dominates all other shocks within the window. Our results suggest that empirical work using the high-frequency event study regression approach should be very careful to isolate the surprise component of the policy news so that the policy shock dominates the other shocks in the event window. Hence, the proposed research provides theoretical support to the concerns expressed recently by Bauer and Swanson (2023a, 2023b) on the credibility of some high-frequency event study estimates when the outcome variable is the Blue Chip forecast revision that is based on a one-month window. Given that the narrow size of the window is not a sufficient condition for identification, empirical researchers should devise creative ways to extract the unexpected component of the news.
Note that the high-frequency event study method is different from heteroskedasticity-based identification [cf. rigobon:2003], sometimes referred to as Rigobon’s method. The latter also uses information from the control sample, which includes windows (either 30-minutes or 1-day windows) that do not bracket an announcement or event. The main heteroskedasticity-based identification condition is that the volatility of the policy shock is larger in the event windows than in the control windows. This is different from relative exogeneity for two reasons. First, the required increase in volatility of the policy shock is not relative to the other shocks in the event window but relative to the policy shock in the control windows. Second, the increase in volatility need not be infinite. Thus, heteroskedasticity-based identification requires neither stronger nor weaker conditions than the relative exogeneity condition. In addition, Rigobon’s estimator is an instrumental variables estimator and therefore different from the OLS high-frequency event study estimator. Since these two estimators correspond to different estimands, they identify different causal effect in general.
@@ 2.4 Relation of the Theoretical Results with the Empirical Literature
There are essentially no formal identification results about high-frequency event studies in the literature. Empirical works have often mentioned that changes in the policy variable in the event windows are dominated by the information about future monetary policy contained in the FOMC announcements [see, e.g., Nakamura and Steinsson (2018a)]. However, precise conditions have not been provided. Rigobon and Sack (2004) noted that the bias of the OLS event study estimator disappears in a stylized linear model when relative exogeneity holds. However, the model they analyzed is admittedly a “clear oversimplification”. As such, this stylized model does not enable practically-relevant identification analysis or econometric results. In addition, showing that the bias disappears in this context does not imply identification of a causal estimand. Nevertheless, we credit Rigobon and Sack’s heuristic analysis as an early insight into understanding the validity of the OLS event study estimator. Gürkaynak and Wright (2013) surveyed the empirical literature and presented useful discussions. They also argued that the event study regression works even when the variance of the policy shock is not large relative to the variance of the other shocks because in a narrow window around an FOMC announcement the policy news can depend on lagged changes in asset returns but not on contemporaneous changes. That is, they wrote $D_{t}=\alpha Y_{t-j}+Z^{\prime}_{t-j}\gamma_{2}+e_{t}$, where $j\geq 1$. However, as we explain in Example 1, the choice of a narrow window is not sufficient for ruling out omitted variables bias. Further, when $Y_{t}$ is the change in an asset price and $D_{t}$ is, for example, the price change in the Federal funds or eurodollar futures, assuming that $Y_{t}$ does not contemporaneously affect $D_{t}$ constitutes a strong empirical restriction.
@@
In recent work, Bauer and Swanson (2023a, 2023b), Cieslak (2018) and Miranda-Agrippino and Ricco (2021) provided empirical evidence for some predictability of $D_{t}$ with publicly available macroeconomic or financial market information that predates the FOMC announcement. This implies that $D_{t}$ does not correctly isolate the unexpected component of the policy surprise and may be in fact simultaneously determined with the outcome variable $Y_{t}$. [^11] This simultaneity follows from the plausible correlation between the publicly available macroeconomic and financial market information prior to the announcement and the macroeconomic factors at the time of the announcement that may affect $Y_{t}$. Bauer and Swanson (2023a) proposed to take the residuals from a regression of those surprises on the economic and financial variables that predate the announcements. So in Bauer and Swanson (2023a) $D_{t}$ is actually the orthogonalized monetary policy surprise rather than the surprise itself. However, Bauer and Swanson (2023a) showed that the orthogonalized policy shocks yield the same results as the unadjusted policy shocks when $Y_{t}$ is measured as the change in a 30-minute window, as for the case of asset prices or Treasury yields. This corroborates our result that some endogeneity of $D_{t}$ does not preclude the validity of the approach if the variance of the unadjusted policy shocks is much larger than that of the other variables in the system.
@@
Evidence of nonlinearities is often documented in the empirical literature [cf. Bauer and Swanson (2023a)]. Our results establish the causal meaning of the event study estimand when the relationship between the outcome and the policy variable is potentially nonlinear as long as the additive separability condition holds.
@@
The leakage documented in the literature does not necessarily imply that relative exogeneity is violated. Intuitively, as long as the key news is revealed with the actual announcements, the high-frequency event study is still characterized by the lumpy manner with which a disproportionate amount of information is unveiled to the public. [^13] This argument applies to the effects of President Trump’s tweets that criticize the Federal Reserve on financial markets documented by Bianchi, Kind, and Kung (2020). They showed that those tweets had a negative effect on the expected Fed funds rate with the magnitude growing by horizon. If the tweet occurs in between the time the survey is collected and the scheduled macroeconomic announcement, then $D_{t}$ does not isolate the expected component of the news correctly. This challenges relative exogeneity. However, if the key macroeconomic data information is revealed in the announcement, then $D_{t}$ is primarily driven by the infinite variance policy shock $e_{t}$ and so missing the effect of the tweets on the updates of the expectations is negligible.
@@
With regards to the choice of the length of the event window, the early literature commonly used a 1-day window while the more recent literature recommended to use narrower windows with the goal of reducing the background noise. The latter refers to variation in $u_{t}$ on the day of the announcement that is related to news other than the policy news. In addition, it is common in the literature to use the same window for both $Y_{t}$ and $D_{t}$. Nakamura and Steinsson (2018a) is an exception as they used a 1-day window for $Y_{t}$ and a 30-minute window for $D_{t}$.
@@
Although a narrower window is associated with a smaller probability of including news about events other than the policy announcement, it also has disadvantages relative to using a longer window for $Y_{t}$. First, with some information leakage, asset prices $Y_{t}$ may respond before the announcement is actually made. Second, with learning or sluggish market adjustments, asset prices may take some time to incorporate the news and so changes in $Y_{t}$ may occur also in the hours after the announcement. [^14] Note that the complications arising from the effect of information leakage and learning on $Y_{t}$ is different from those we discuss in Section 2.5 for proper construction of $D_{t}$. Intuitively, the impact of information leakage on $Y_{t}$ depends on the potential causal effect of the policy news on $Y_{t}$. In the linear model (2.1) this is captured by $\beta$. If $\beta=0$, then information leakage has no effect on $Y_{t}$ while it does complicate the proper construction of the surprise $D_{t}$ irrespective of the value of $\beta$. Our analysis shows that the key identification condition (relative exogeneity) does not refer to the size of the window, it only requires that, whatever window length is chosen, the policy shock dominates any other shock within that window.
@@
Theorem 1 allows us to provide a formal explanation for a recent debate on some puzzling event study regression results documented in the literature. It has been shown that regressions of private-sector macroeconomic forecast revisions on monetary policy surprises often produce coefficients with signs opposite to those of standard macroeconomic models. [^15] For example, a surprise monetary policy tightening is associated with a statistically significant upward revision in the Blue Chip consensus forecasts for real GDP growth. This is inconsistent with the standard macroeconomic view that a monetary policy tightening should cause future GDP to fall. Campbell, Evans, Fisher, and Justiniano (2012), Nakamura and Steinsson (2018a) and Romer and Romer (2000) argued in favor of the “Fed information effect” for which these puzzling results are due to monetary policy surprises revealing private information held by the Federal Reserve. Bauer and Swanson (2023a) challenged these views, arguing that these event study estimates suffer from omitted variables bias. In Section 4, we analyze this puzzle in detail and show that in the event study regressions involving the Blue Chip forecasts, it is unlikely that relative exogeneity holds because the forecasts are evaluated at a much lower frequency than the policy variable. Intuitively, while the policy surprise $D_{t}$ is constructed as a 30-minute change, $Y_{t}$ is the one-month change in the Blue Chip forecasts and so the latter likely have a large variance relative to the former as they aggregate all news and factors that are relevant over the month. Thus, it becomes important to control for macroeconomic and financial variables that predate the announcements. This explains why using the orthogonalized shocks indeed allowed Bauer and Swanson (2023a) to overturn the documented puzzling estimates. In contrast, when $Y_{t}$ is a 30-minute or 1-day change in an asset price or Treasury yield, the variance of $D_{t}$ is much larger in relative terms and can eliminate the endogeneity arising from omitted variables that predate the FOMC announcement.

### v4

@@ 2.2.1 Potential Outcomes Framework
We assume that $d\in\mathbf{D}$ for an appropriate set $\mathbf{D}$. A potential outcome $Y_{t}\left(d\right)$ describes which value the outcome would have taken at time $t$ under treatment value $d$. The definition implies that the potential outcome $Y_{t}\left(d\right)$ does not depend on future treatments. Rambachan and Shephard (2021) described this property as “non-anticipating potential outcomes”. The definition also implies that the potential outcome $Y_{t}\left(d\right)$ does not depend on past treatments (unless they are elements of $Z_{t}$). This is realistic in the high-frequency event study setting for two reasons. First, $Y_{t}$ typically measures a change in an asset price or a survey forecast within a narrow window around a policy announcement. Thus, market efficiency or rational expectations, respectively, imply that market participants or professional forecasters use all public information available at the start of the window and so $Y_{t}$ is constructed by conditioning on that information. Second, for $t\in\mathbf{P}$ the latest past treatment is $D_{t-1}$ ($(\{t-1\}\in\mathbf{C})$ which captures the surprise change in a policy within a 30-minute window that does not involve any policy announcement and so the change is very small or equal zero. The potential outcome should not be confused with the outcome $Y_{t}=Y_{t}\left(D_{t}\right)$. Finally, note that in terms of the structural function in (2.3), $Y_{t}(d)=\varphi_{Y}(d\,,Z_{t}\,,u_{t}\,,t)$.
@@
Under these assumptions, the MCE of the time $t$ policy on the time $t$ normalized potential outcome is defined as $\partial\widetilde{Y}_{t}\left(d\right)/\partial d$. Assumptions 5-6 involve the normalized quantities instead of the actual quantities since we analyze estimands and estimators under relative exogeneity (2.4), for which the support of $e_{t}$ and $D_{t}$ necessarily become unbounded. Following Rambachan and Shephard (2021), to apply basic tools such as the fundamental theorem of calculus we need the argument of the relevant function to have support on a closed interval.
@@ 2.3 Identification Results
where $H_{t}\left(d\right)=\mathbf{1}\{d\leq\widetilde{D}_{t}\}(\widetilde{D}_{t}-\mathbb{E}(\widetilde{D}_{t}))$ with $\mathbb{E}(H_{t}\left(d\right))\geq 0,$ and $\int_{\mathbf{D}}\mathbb{E}\left(H_{t}\left(d\right)\right)\mathrm{d}d=1$,
$\Delta_{t}=\mathbb{E}\left[\widetilde{Y}_{t}\left(\underline{d}\right)\left(\widetilde{D}_{t}-\mathbb{E}\left(\widetilde{D}_{t}\right)\right)\right]=\mathrm{Cov}\left(\widetilde{Y}_{t}\left(\underline{d}\right),\,\widetilde{D}_{t}\right),$
and $\left|\Delta_{t}\right|$ is decreasing in $\sigma_{e,t}^{2}$.
@@
For the weighted average of the MCEs in Theorem 2, a higher weight
$\mathbb{E}\left(H_{t}\left(d\right)\right)$ $=\mathbb{E}\left(\widetilde{D}_{t}-\mathbb{E}(\widetilde{D}_{t})|\,d\leq\widetilde{D}_{t}\right)\times\mathbb{P}\left(d\leq\widetilde{D}_{t}\right)$
is not necessarily given to large values of $d$ because large values of $d$ may be associated with small tail probabilities of the distribution of $\widetilde{D}_{t}$, $\mathbb{P}(d\leq\widetilde{D}_{t})$. The intuition on how $\beta_{\mathrm{ES},t}$ is able to recover a weighted average of MCEs under relative exogeneity is as follows. Relative exogeneity, in combination with Assumptions 1-3, imply that when appropriately normalized, the potential outcome $Y_{t}\left(d\right)$ and the policy variable $D_{t}$ behave as if they are uncorrelated. Intuitively, the effect of the policy shock on $D_{t}$ is not influenced by $Y_{t}$ (separability) and so changes in the policy are entirely determined by changes in the policy shock (relative exogeneity). In addition, as $\sigma_{e,t}^{2}\rightarrow\infty$ the relative bias generated by the omitted variables $Z_{t}$ becomes negligible since the correlation between $Z_{t}$ and $D_{t}$ is an order of magnitude smaller than the variation in $D_{t}$ generated by $e_{t}$.
@@
The key identification condition is relative exogeneity, which also relates to the size of the time window surrounding an announcement or event. As the size of the window expands, it becomes less likely that the policy shock dominates all other variables within the window. Given that the narrow size of the window is not a sufficient condition for identification, our results suggest that empirical work using the high-frequency event study regression approach should be very careful to isolate the surprise component of the policy news so that the policy shock dominates the other variables in the event window. Hence, our theoretical results support the concerns expressed recently by Bauer and Swanson (2023a, 2023b) on the credibility of some high-frequency event study estimates when the outcome variable is the Blue Chip forecast revision that is based on a one-month window.
Note that the high-frequency event study method is different from heteroskedasticity-based identification [cf. rigobon:2003], sometimes referred to as Rigobon’s method. The latter also uses information from the control sample, which includes windows (either 30-minute or 1-day windows) that do not bracket an announcement or event. The main heteroskedasticity-based identification condition is that the volatility of the policy shock is larger in the event windows than in the control windows. This is different from relative exogeneity for two reasons. First, the required increase in volatility of the policy shock is not relative to the other variables in the event window but relative to the policy shock in the control windows. Second, the increase in volatility need not be infinite. Thus, heteroskedasticity-based identification requires neither stronger nor weaker conditions than the relative exogeneity condition. In addition, Rigobon’s estimator is an instrumental variables estimator and therefore different from the OLS high-frequency event study estimator. Since these two estimators correspond to different estimands, they identify different causal effect in general.
@@ 2.4 Relation of the Theoretical Results with the Empirical Literature
There are essentially no formal identification results about high-frequency event studies in the literature. Empirical works have often mentioned that changes in the policy variable in the event windows are dominated by the information about future monetary policy contained in the FOMC announcements [see, e.g., Nakamura and Steinsson (2018a)]. However, precise conditions have not been provided. Rigobon and Sack (2004) noted that the bias of the OLS event study estimator disappears in a stylized linear model when relative exogeneity holds. However, the model they analyzed is admittedly a “clear oversimplification”. As such, this stylized model does not enable practically-relevant identification analysis or econometric results. In addition, showing that the bias disappears in this context does not imply identification of a causal estimand. Nevertheless, we credit Rigobon and Sack’s heuristic analysis as an early insight into understanding the validity of the OLS event study estimator. Gürkaynak and Wright (2013) surveyed the empirical literature and presented useful discussions. They also argued that the event study regression works even when the variance of the policy shock is not large relative to the variance of the other variables because in a narrow window around an FOMC announcement the policy news can depend on lagged changes in asset returns but not on contemporaneous changes. That is, they wrote $D_{t}=\alpha Y_{t-j}+Z^{\prime}_{t-j}\gamma_{2}+e_{t}$, where $j\geq 1$. However, as we explain in Example 1, the choice of a narrow window is not sufficient for ruling out omitted variables bias. Further, when $Y_{t}$ is the change in an asset price and $D_{t}$ is, for example, the price change in the Federal funds or eurodollar futures, assuming that $Y_{t}$ does not contemporaneously affect $D_{t}$ constitutes a strong empirical restriction.
@@
In recent work, Bauer and Swanson (2023a, 2023b), Cieslak (2018) and Miranda-Agrippino and Ricco (2021) provided empirical evidence for some predictability of $D_{t}$ with publicly available macroeconomic or financial market information that predates the FOMC announcement. This implies that $D_{t}$ does not correctly isolate the unexpected component of the policy surprise and may be in fact simultaneously determined with the outcome variable $Y_{t}$. [^11] This simultaneity follows from the plausible correlation between the publicly available macroeconomic and financial market information prior to the announcement and the macroeconomic factors at the time of the announcement that may affect $Y_{t}$. Bauer and Swanson (2023a) proposed to take the residuals from a regression of those surprises on the economic and financial variables that predate the announcements. So in Bauer and Swanson (2023a) $D_{t}$ is actually the orthogonalized monetary policy surprise rather than the surprise itself. However, Bauer and Swanson (2023a) showed that the orthogonalized policy variables yield the same results as the unadjusted policy variables when $Y_{t}$ is measured as the change in a 30-minute window, as for the case of asset prices or Treasury yields. This corroborates our result that some endogeneity of $D_{t}$ does not preclude the validity of the approach if the variance of the unadjusted policy shocks is much larger than that of the other variables in the system.
@@
Evidence of nonlinearities is often documented in the empirical literature [cf. Bauer and Swanson (2023a)]. Our results establish the causal meaning of the event study estimand when the relationship between the outcome and the policy variable is potentially nonlinear as long as the additive separability conditions hold.
@@
As we discuss below, the leakage documented in the literature does not necessarily imply that relative exogeneity is violated. Intuitively, as long as the key news is revealed with the actual announcements, the high-frequency event study is still characterized by the lumpy manner with which a disproportionate amount of information is unveiled to the public. [^13] This argument applies to the effects of President Trump’s tweets that criticize the Federal Reserve on financial markets documented by Bianchi, Kind, and Kung (2020). They showed that those tweets had a negative effect on the expected Fed funds rate with the magnitude growing by horizon. If the tweet occurs in between the time the survey is collected and the scheduled macroeconomic announcement, then $D_{t}$ does not isolate the expected component of the news correctly. This challenges relative exogeneity. However, if the key macroeconomic data information is revealed in the announcement, then $D_{t}$ is primarily driven by the infinite variance policy shock $e_{t}$ and so missing the effect of the tweets on the updates of the expectations is negligible.
@@
With regards to the choice of the length of the event window, the early literature commonly used a 1-day window while the more recent literature recommended to use narrower windows with the goal of reducing the background noise. In addition, it is common in the literature to use the same window for both $Y_{t}$ and $D_{t}$. Nakamura and Steinsson (2018a) is an exception as they used a 1-day window for $Y_{t}$ and a 30-minute window for $D_{t}$.
@@
Although a narrower window is associated with a smaller probability of including news about events other than the policy announcement, it also has disadvantages relative to using a longer window for $Y_{t}$. First, with some information leakage, asset prices $Y_{t}$ may respond before the announcement is actually made. Second, with learning or sluggish market adjustments, asset prices may take some time to incorporate the news and so changes in $Y_{t}$ may occur also in the hours after the announcement. [^14] Note that the complications arising from the effect of information leakage and learning on $Y_{t}$ are different from those we discuss in Section 2.5 for proper construction of $D_{t}$. Intuitively, the impact of information leakage on $Y_{t}$ depends on the potential causal effect of the policy news on $Y_{t}$. In the linear model (2.1) this is captured by $\beta$. If $\beta=0$, then information leakage has no effect on $Y_{t}$ while it does complicate the proper construction of the surprise $D_{t}$ irrespective of the value of $\beta$. Our analysis shows that the key identification condition (relative exogeneity) does not explicitly refer to the size of the window, it only requires that, whatever window length is chosen, the policy shock dominates any other shock within that window.
@@
Theorem 1 allows us to provide a formal explanation for a recent debate on some puzzling event study regression results documented in the literature. It has been shown that regressions of private-sector macroeconomic forecast revisions on monetary policy surprises often produce coefficients with signs opposite to those of standard macroeconomic models. [^15] For example, a surprise monetary policy tightening is associated with a statistically significant upward revision in the Blue Chip consensus forecasts for real GDP growth. This is inconsistent with the standard macroeconomic view that a monetary policy tightening should cause future GDP to fall. Campbell, Evans, Fisher, and Justiniano (2012), Nakamura and Steinsson (2018a) and Romer and Romer (2000) argued in favor of the “Fed information effect” for which these puzzling results are due to monetary policy surprises revealing private information held by the Federal Reserve. Bauer and Swanson (2023a) challenged these views, arguing that these event study estimates suffer from omitted variables bias. In Section 4, we analyze this puzzle in detail and show that in the event study regressions involving the Blue Chip forecasts, it is unlikely that relative exogeneity holds because the forecast revisions are evaluated at a much lower frequency than the policy variable. Intuitively, while the policy surprise $D_{t}$ is constructed as a 30-minute change, $Y_{t}$ is the one-month change in the Blue Chip forecasts and so the latter likely has a large variance relative to the former as it aggregates all news and factors that are relevant over the month. Thus, it becomes important to control for macroeconomic and financial variables that predate the announcements in this context. This explains why using the orthogonalized shocks indeed allowed Bauer and Swanson (2023a) to overturn the documented puzzling estimates. In contrast, when $Y_{t}$ is a 30-minute or 1-day change in an asset price or Treasury yield, the variance of $D_{t}$ is much larger in relative terms and can eliminate the endogeneity arising from omitted variables that predate the FOMC announcement.